# Instabilities of Gauged Q-Balls

### v1

@@ VI Discussion and Conclusion
We have also used our formula to characterize the boundary between stable and unstable parameters. This boundary does not agree perfectly with the results of Kinach:2022jdx . It appears that any Q-ball found to be unstable by our formula is also found to be unstable in the numerical calculations; however, there appear to be Q-balls which are found numerically to be unstable which are not captured by our analysis. This suggests strongly that there are other perturbations that we have missed that will extend the region of instability. On a related note, the region of instability in the numerical results appears to be linked to the condition ${d\omega\over dQ}>0$, at least for small gauge couplings. We do not see such a link, which is again suggestive that a different form of perturbation may be relevant for Q-balls. This analysis is left for future work.
@@ VII Acknowledgments
This work was supported in part by NSF Grant No. PHY-2210283.

### v3

@@ VI Discussion and Conclusion
Finally, we note that our predicted boundary between stable and unstable parameters does not agree perfectly with the results of Kinach:2022jdx . It appears that any Q-ball found to be unstable by our formula is also found to be unstable in the numerical calculations; however, there appear to be Q-balls which are found numerically to be unstable which are not captured by our analysis. This suggests strongly that there are other perturbations that we have missed that will extend the region of instability. On a related note, the region of instability in the numerical results appears to be linked to the condition ${d\omega\over dQ}>0$, at least for small gauge couplings. We do not see such a link, which is again suggestive that a different form of perturbation may be relevant for Q-balls. This analysis is left for future work.
@@ VII Acknowledgments
This work was supported in part by NSF Grant No. PHY-2210283. We thank M. Smolyakov for a comment on an initial version of this paper.